• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Deur

    Log entry time 22:19:59 on July26,2001

    Entry number 67841

    keyword=Epics problem

    We have an epics problem (a min after that the beam was back!). The target diplays, the charge asy feedback, the raster mcc information and the spectrometer position info are blanck. the laser absorption spectrum is flat.
    The MEDM message that appeared said that it is a network problem.
    Reboot of the target ioc (iocha26) did not work. We cannot connect or ping it either.
    Kathy was paged and she is investigating.